OpenStudy (anonymous):

Chloe Closson paid for her vacation to the mountains with a $3,000 installment loan at 10% for 12 months. Her monthly payments are $263.75. After five payments, the balance due was $1,786.20. If Chloe pays off the loan with the sixth payment, how much will she save? Round up to the nearest cent.

OpenStudy (agent0smith):

Hmm, i guess we can just work out how much interest she would pay over the remaining... 7 months? and then use that to find the savings

OpenStudy (agent0smith):

I'm not 100% sure on these questions though, because they're finance-based. I think we just need to add up all the remaining payments, and compare it to the 1786.20 she'd pay now...

OpenStudy (anonymous):

Yeah, I'm not getting these either. @agent0smith

OpenStudy (agent0smith):

Well let's just see how much would be left...after 5 payments she owes 1786.20, let's assume that's what she pays on the 6th payment... idk if we have to account for the interest, but if you do we can change the answer later, maybe If she made 7 payments of 263.75: that adds up to 1846.25 so she'd have saved 1846.25-1786.20 = $60.05

OpenStudy (anonymous):

I feel like this question is insolvable. I have literally seen like 10 answers for this question and every singe one has been wrong. @agent0smith

OpenStudy (agent0smith):

Just in case, i'll account for the interest... the interest for one month at 10% is 0.1/12 (12 months) then use the compound interest formula: A = 1786.20(1+0.1/12) = 1801.085 so now the difference would be 1846.25 - 1801.085 = $45.16 saved

OpenStudy (agent0smith):

The payment seems small... 263.75 for 12 payments is only $3165... i'd think a 10% loan would have AT least $3300 in total payments, since you'd expect there to be at least 10% interest which is $300.

OpenStudy (anonymous):

Haha that still wasn't it. But it's okay. Thank you though! :) @agent0smith

OpenStudy (agent0smith):

Yeah the numbers in the question seem a little off. I couldn't figure out why the amount saved seemed so small, relative to the interest each month.

OpenStudy (agent0smith):

But maybe it's just cos i'm missing some financial algebra knowledge that's required...